OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7> -8.3° 1.7° 3.3° 13.3°

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

first you need to compute the dot product

jimthompson5910 (jim_thompson5910):

what do you get when you do that

OpenStudy (anonymous):

kk onesec

OpenStudy (anonymous):

{72, 49}

jimthompson5910 (jim_thompson5910):

add those two up

OpenStudy (anonymous):

121

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (cwrw238):

now you need to find the modulus of each vector

OpenStudy (anonymous):

how do u do that

OpenStudy (cwrw238):

mod of u = sqrt(8^2 + 7^2)

OpenStudy (cwrw238):

mod of v = sqrt(9^2 + 7^2)

OpenStudy (anonymous):

sqrt(8^2 + 7^2)=113^(1/2) sqrt(9^2 + 7^2)=130^(1/2)

OpenStudy (cwrw238):

ok now cos x = dot product / ( mod u * mod v) where x is the required angle

OpenStudy (cwrw238):

what do you get for modu * mod v ?

OpenStudy (anonymous):

121 / ( 113^(1/2) *130^(1/2)) ???

OpenStudy (anonymous):

modu * mod v = Sqrt[14690]

OpenStudy (cwrw238):

no use your calculator to work out the 2 square roots seperately then multiply them

OpenStudy (anonymous):

121.202

OpenStudy (cwrw238):

sqrt 113 = 10.63015 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

modu * mod v =121

OpenStudy (cwrw238):

ok so we have cos x = 121 / 121.202

OpenStudy (anonymous):

cos(x) = 0.998333

OpenStudy (cwrw238):

right now use your calculator to find x

OpenStudy (cwrw238):

its the cos-1 key or ( on some its arcos)

OpenStudy (anonymous):

3.3

OpenStudy (cwrw238):

yep

OpenStudy (cwrw238):

the important formula to remember is u.v = |u||v| cos x where u and v are the vectors and |u| and |v| are the modulus of u and v

OpenStudy (cwrw238):

@jim_thompson5910 is it modulus or magnitude the correct term?

jimthompson5910 (jim_thompson5910):

they are the same thing pretty much

jimthompson5910 (jim_thompson5910):

the modulus of a vector = the magnitude of a vector

OpenStudy (cwrw238):

oh ok